Two spherical objects have masses of 250 kg and 550 kg. Their centers are separated by a distance of 45 m. Find the gravitational attraction between them.

1
Expert's answer
2021-11-18T15:14:13-0500

"m_1=250kg\\m_2=550kg\\d=45m"

"F=\frac{Gm_1m_2}{r^2}"


"F=\frac{6.67\times10^{-11}\times250\times550}{45^2}=4.52\times10^{-9}N"